a rectangle perimeter is 78 inches
the length is 6 more than twice its width.
what is the length and width?

Answer by mananth(16946)

let width be x
length = 2x+6
...
perimeter = 78
2*(l+w)=perimeter
2*(x+(2x+6))=78
2(3x+6)=78
6x+12=78
-12
6x=78-12
6x=66
x= 11 the width
length = 2x+6 = 22+6 = 28
\..
CHECK 28+28+11+11= 78
...
m.ananth@hotmail.ca